OpenStudy (aravindg):

ax/cos theta)+(by/sin theta)=a^2-b^2 and (axsin theta/cos^2 theta)-(bycos theta/sin^2 theta)=0 show that (ax)^2/3 +(by)^2/3=(a^2-b^2)^2/3

OpenStudy (anonymous):

Use Equation editor...

OpenStudy (aravindg):

help

OpenStudy (anonymous):

\[\frac{ax}{cos\theta} + \frac{by}{sin\theta} = a^2 -b^2\]\[\frac{ax sin \theta}{cos^2 \theta }-\frac{bycos\theta}{sin^2\theta} = 0\]

OpenStudy (aravindg):

\[ax^{2/3}+by ^{2/3}=( a ^{2}-b ^{2})^{2/3}\]

OpenStudy (aravindg):

helpppppp

OpenStudy (anonymous):

is it not (ax)^2/3 +(by)^2/3=... ?

OpenStudy (aravindg):

no it is to prove